A man and his brother have a combined age of 162 years. 49 years ago, the man was 7 times older than his brother. How old is the man? How old is the brother?
Answer by edjones(7311)

Let m=the man's age and b=brother's
m+b=162
m-49=7(b-49)
.
m=162-b
162-b-49=7b-343
8b=343+162-49
8b=456
b=57 years old
m=162-57=105 yrs old
.
Ed